Two spherical balls each of mass 1kg are placed 1cm apart. The gravitational force of attraction between them in Newton is

Grade:6

3 Answers

Arun
5 years ago
F = GMm / r^2
= 6.67 * 10^-11 × 1 × 1 / 1^2
= 6.67 × 10^-11 N

Gravitational force between them is 6.67 × 10^-11 N
Khimraj
5 years ago
Gravitational force between two body of mass m and M seprated by distance r is given by
F = GMm / r2
= 6.67 * 10-11 × 1 × 1 / 12
= 6.67 × 10-11 N
Baby harshitha
4 years ago
F=Gm1m2/r^2
Given,m1=m2=1 kg 
r=1cm=10^-2meters
Therefore G=6.67×10^-11N-m^2/kg^2
F=6.67×10^-11×1×1/(10^-2)^2
6.67×10^-11/10^-4 = 6.67×10^-11+4 = 6.67×10^-7
F=6.67×10^-7